Home‎ > ‎Crew 131 - Embry Riddle‎ > ‎

1209-EngineeringReport

Engineering Report - Evening
12/09/2013
Chui Lee
Was Power Interrupted:
no
Time Of Power Outage:

Length Of Power Outage:

Diesel Fuel Level:
.5
Propane Fuel Level:
.3
Gasoline Fuel Level:
0.0
Fuel Notes And Comments:

Atv Oil Quarts:
0
Car Oil Quarts:
3.5
Oil Notes And Comments:

Trailer Potable Water Level:
14.224
Trailer Potable Water Refilled:
no
Outside Potable Water Level:
45.44
Outside Potable Water Refilled:
no
Loft Potable Water Gallons:
45
Loft Potable Water Refilled:
yes
Potable Water Meter Reading:
706223
Greenhouse Notes And Comments:
Heater/pilot light is on at 75 degrees
Telescope Used:
no
Observatory Notes And Comments:

Network Failed:
no
Network Throttled:
no
Network Signal Strength:
192 (13.92dB)
Network Download Allowane:
29
Network Notes And Comments:

Hvac Air Conditioner Used:
no
Hvac Furnace Used:
no
Hvac Lab Wall Heater Used:
no
Hvac Upper Deck Thermostat Setting:
75
Hvac Upper Deck Temperature Reading:
72
Hvac Lower Deck Thermostat Setting:
75
Hvac Lower Deck Temperature Reading:
53.6
Hvac Notes And Comments:

Hab Car Used:
no
Hab Car Oil Checked:
no
Hab Car Tire Status:
Good
Hab Car Odometer:
130658.3
Hab Car Notes And Comments:

Atv One Used:
no
Atv One Oil Checked:
yes
Atv One Fuel Used Gals:
0
Atv One Tires Status:
Good
Atv One Hours Used Day:
0
Atv One Hours Used Year:
0
Atv One Notes And Comments:

Atv Two Used:
no
Atv Two Oil Checked:
yes
Atv Two Fuel Used Gals:
0
Atv Two Tires Status:
Good
Atv Two Hours Used Day:
0
Atv Two Hours Used Year:
.083
Atv Two Notes And Comments:

Atv Three Used:
no
Atv Three Oil Checked:
yes
Atv Three Fuel Used Gals:
0
Atv Three Tires Status:
Good
Atv Three Hours Used Day:
0
Atv Three Hours Used Year:
0
Atv Three Notes And Comments:

Atv Four Used:
no
Atv Four Oil Checked:
yes
Atv Four Fuel Used Gals:
0
Atv Four Tires Status:
Good 
Atv Four Hours Used Day:
0
Atv Four Hours Used Year:
0
Atv Four Notes And Comments:

Atv Five Used:
no
Atv Five Oil Checked:
no
Atv Five Fuel Used Gals:

Atv Five Tires Status:

Atv Five Hours Used Day:

Atv Five Hours Used Year:

Atv Five Notes And Comments:

Comments:
>>> Static tank (before filling): 45.72
>>> Static tank (after filling): 45.44
>>> Loft Tank (before filling): 20
>>> Loft Tank (after filling): 45
>>> Trailer tank (before filling): 14.224
>>> Trailer tank (after filling): 14.224



>> Pilot light/heater in Green Hab is lit/heated to 70-75.

>> Upon use of the EVA suits, it appears the fittings have been switched between all the suits and helmets. As instructed by Judd, I will be switching the tubes to their correct suits.

>> As instructed by Judd-- I opened up a section for the controls of Pack 8, and reinforced it with the protective foam tubing (see attached photo in Mission Control e-mail). To try and increase the air flow/performance of the packs, we tried opening the bottom of Pack 9 (see attached photo in Mission Control e-mail). We left the excess cloth in case this was not the optimal change. Currently allow Pack 9 to charge, to test tomorrow. Will send updates.

_________________________________

Engineering Report
12/09/2013 - Morning Report
Chui Lee


Was Power Interrupted:
no
Time Of Power Outage:

Length Of Power Outage:

Diesel Fuel Level:
0.5
Propane Fuel Level:
.3
Gasoline Fuel Level:
0.0
Fuel Notes And Comments:

Atv Oil Quarts:
0
Car Oil Quarts:
3.25
Oil Notes And Comments:

Trailer Potable Water Level:
14.224
Trailer Potable Water Refilled:
no
Outside Potable Water Level:
45.72
Outside Potable Water Refilled:
no
Loft Potable Water Gallons:
45
Loft Potable Water Refilled:
no
Potable Water Meter Reading:
706161.0
Greenhouse Notes And Comments:
Still attempting to turn on pilot lights and heater. Found the pilot light and gas valve.
Telescope Used:
no
Observatory Notes And Comments:

Network Failed:
no
Network Throttled:
no
Network Signal Strength:
189 (13.87dB)
Network Download Allowane:
46%
Network Notes And Comments:
Currently keeping the HAB computer off. Significant decrease in bandwidth use.
Hvac Air Conditioner Used:
no
Hvac Furnace Used:
yes
Hvac Lab Wall Heater Used:
no
Hvac Upper Deck Thermostat Setting:
75
Hvac Upper Deck Temperature Reading:
72
Hvac Lower Deck Thermostat Setting:
75
Hvac Lower Deck Temperature Reading:
53.6
Hvac Notes And Comments:

Hab Car Used:
no
Hab Car Oil Checked:
no
Hab Car Tire Status:
Good
Hab Car Odometer:
130658.3
Hab Car Notes And Comments:

Atv One Used:
no
Atv One Oil Checked:
no
Atv One Fuel Used Gals:
0
Atv One Tires Status:
Good
Atv One Hours Used Day:
0
Atv One Hours Used Year:
0
Atv One Notes And Comments:

Atv Two Used:
no
Atv Two Oil Checked:
no
Atv Two Fuel Used Gals:
0
Atv Two Tires Status:
Good
Atv Two Hours Used Day:
0
Atv Two Hours Used Year:
.083
Atv Two Notes And Comments:

Atv Three Used:
no
Atv Three Oil Checked:
no
Atv Three Fuel Used Gals:
0
Atv Three Tires Status:
Good
Atv Three Hours Used Day:
0
Atv Three Hours Used Year:
0
Atv Three Notes And Comments:

Atv Four Used:
no
Atv Four Oil Checked:
no
Atv Four Fuel Used Gals:
0
Atv Four Tires Status:
Good
Atv Four Hours Used Day:
0
Atv Four Hours Used Year:
0
Atv Four Notes And Comments:

Atv Five Used:
no
Atv Five Oil Checked:
no
Atv Five Fuel Used Gals:

Atv Five Tires Status:

Atv Five Hours Used Day:

Atv Five Hours Used Year:

Atv Five Notes And Comments:

Comments:
MORNING REPORT:
>>> Static tank (before filling): 46.228
>>> Static tank (after filling): 45.72
>>> Loft Tank (before filling): 10
>>> Loft Tank (after filling): 45

>> The water levels of the static tank fluctuates due to the extreme temperatures, it tends to freeze overnight and through the day. Currently turning on pump every 2 hours to ensure pipes do not freeze.
>> DG told us he will be bringing the tank to refill tomorrow!



>>> As per Mission Control's request, we are currently in the process of trying to light the pilot lights/turn on the heat in the Green Hab. Will submit updates later. 

>>> EVA Suits Update:
#1) Fan is working with optimal performance, the helmet has undergone sufficient repairs near the front face and around the neck attachment. Approved for use.
#2) Fan is not working in optimal performance (incredibly weak). When opened up, the filters seemed to be quite clogged with dust and particulates. Attempted to clear filters but did not conclude with a significant increase in performance. Denied for use. 
#3) Fan is working with optimal performance. The entire right side of the helmet has undergone extensive repairs and is almost fully covered in tape due to cracking. Approved for use.
#4) Fan is not working in optimal performance (weak, but not as significantly as #2). We are assuming it is struggling with the same issues with the filters. The left air tube has had significant repair. Helmet attachment on the EVA pack is currently missing one of it's straps. Helmet face has crack that is not fixed. Not approved for use.
#5) Fan is not working in optimal performance (weak). Likely attributed to filters. Helmet has minimal damage. Not approved for use.
#6) Fan is working with optimal performance. Helmet has long cracked across the face, not currently fixed as it will remove visibility. Approved for use.

(NEW/EXPERIMENTAL PACKS)- We have not used them as they were not charged for our first EVA today and we were conducting a steep hill EVA. Tomorrow we will be conducting an ATV EVA, thus we will not likely use the suits.
#7) Fan is not currently working, missing hoses. -- Currently charging to ensure it is not a dead battery. Will update.
#8) Fan is not working with optimal performance (weak). Currently charging to ensure it is not the battery as the cause. Missing opening for switches and charger.
#9)Fan is not working with optimal performance (weak). 
#10) Fan is not working with optimal performance (weak).

In comparison to the original packs, the experimental packs seem to lack the strength in the fan performance, air flow is weak. The helmets have not been tested aside from a visual evaluation.

>>> Requesting further actions for repair(s). <<<